Determine whether the following DE is seperable, linear , exact, homogeneous, and Bernoulli. next find its solutions by respective methods.


(1+2y^2)dy = ycosxdx

1
Expert's answer
2021-11-01T19:05:03-0400

Rewrite in the form of a first order separable ODE


"\dfrac{1+2y^2}{y}dy=\cos xdx"

Integrate


"\int\dfrac{1+2y^2}{y}dy=\int\cos xdx"

"\ln (|y|)+y^2=\sin x+C"